# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v2

@@ E.1 Computing Resources
E.2 Molecule Generation
@@
We evaluate the quality of generated molecules on two well-known molecular datasets: QM9 (Ramakrishnan et al., 2014) and ZINC250k (Irwin et al., 2012), and obtain the intermediate higher-order skeletons using the 2-cell complex filtering. In experiments, each molecule is preprocessed into a graph comprising adjacency matrix $\bm{A}\in\{0,1,2,3\}^{n\times n}$ and node feature matrix $\bm{X}\in\{0,1\}^{n\times d}$, where $n$ denotes the maximum number of atoms in a molecule of the dataset, and $d$ is the number of possible atom types. The entries of $\bm{A}$ indicate the bond types: 0 for no bound, 1 for the single bond, 2 for the double bond, and 3 for the triple bond. Further, we scale $\bm{A}$ with a constant scale of 3 in order to bound the input of the model in the interval [0, 1], and rescale the final sample of the generation process to recover the bond types. Following the standard procedure (Shi et al., 2020; Luo et al., 2021), all molecules are kekulized by the RDKit library (Landrum et al., 2016) with hydrogen atoms removed. In addition, we make use of the valency correction proposed by Zang & Wang (2020). After generating samples by simulating the reverse diffusion process, the adjacency matrix entries are quantized to discrete values 0,1,2,3{0,1,2,3} by by applying value clipping. Specifically, values in $(-\infty,0.5)$ are mapped to 0, $[0.5,1.5)$ to 1, $[1.5,2.5)$ to 2, and $2.5,+\infty)$ to 3, ensuring the bond types align with their respective categories.
To comprehensively assess the quality of the generated molecules across datasets, we evaluate 10,000 generated samples using several key metrics: validity, validity w/o check, Frechet ChemNet Distance (FCD) (Preuer et al., 2018), Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa & Grave, 2010), uniqueness, and novelty (Jo et al., 2022). FCD quantifies the similarity between generated and test molecules by leveraging the activations of ChemNet’s penultimate layer, assessing the generation quality within the chemical space. In contrast, NSPDK-MMD evaluates the generation quality from the graph topology perspective by computing the MMD between the generated and test sets while considering both node and edge features. Validity is measured as the fraction of valid molecules to all generated molecules after applying post-processing corrections such as valency adjustments or edge resampling, while validity w/o correction, following Jo et al. (2022), computes the fraction of valid molecules before any corrections, providing insight into the intrinsic quality of the generative process. Whether molecules are valid is generally determined by compliance with the valence rules in RDKit (Landrum et al., 2016). Novelty assesses the model’s ability to generalize by calculating the percentage of generated graphs that are not subgraphs of the training set, with two graphs considered identical if isomorphic. Uniqueness quantifies the diversity of generated molecules as the ratio of unique samples to valid samples, removing duplicates that are subgraph-isomorphic, ensuring variety in the output. We report the baseline results taken from Jo et al. (2022) and Kong et al. (2023).
E.3 Generic Graph Generation
@@
To display the topology distribution learning ability, we assess HOG-Diff over four common generic graph datasets: (1) Community-small, containing 100 randomly generated community graphs; (2) Ego-small, comprising 200 small ego graphs derived from the Citeseer network dataset; (3) Enzymes, featuring 587 protein graphs representing tertiary structures of enzymes from the BRENDA database; (4) Stochastic Block Model (SBM), a larger-scale dataset discussed separately in Sec. F.2 due to its distinct evaluation protocol.
@@
Tab. 5 summarizes the key characteristics of the datasets utilized in this study. The table outlines the type of dataset, the total number of graphs, and the range of graph sizes ($|V|$). Additionally, it also provides the number of distinct node types and edge types for each dataset. Notably, the Community-small and Ego-small datasets contain relatively small graphs, whereas Enzymes, SBM and the molecular datasets (QM9 and ZINC250k) exhibit greater diversity in terms of graph size and complexity.
@@
Dataset Graph type #Graphs #Nodes
Node
types
Edge
types
@@ Appendix F Additional Experimental Results
F.1 Analysis of Diffusion Domain Choice
@@
F.2 Scalability Evaluation on Large Graphs
To further evaluate the scalability and robustness of HOG-Diff, we report results on the Stochastic Block Model (SBM) dataset in Tab. 7, which comprises graphs of larger scale and is evaluated under a distinct experimental protocol from that used in Sec. 4.2. The dataset consists of 200 synthetic graphs generated using the stochastic block model. The number of communities is uniformly sampled between 2 and 5, and the number of nodes within each community is uniformly sampled between 20 and 40. Edges are created with probabilities of 0.3 for intra-community connections and 0.05 for inter-community connections.
We adopt the evaluation setting introduced by Martinkus et al. (2022), including the same data splits, feature initialization, and the use of Total Variation (TV) distance to compute the MMDs. The TV kernel is adopted since it offers higher computational efficiency compared to the Earth Mover’s Distance (EMD) kernel, especially for large graphs. The baselines are sourced from (Vignac et al., 2023; Jo et al., 2024) or reproduced using the corresponding publicly available code. As shown in Tab. 7, HOG-Diff achieves competitive performance compared to the state-of-the-art.
Method Deg. $\downarrow$ Clus. $\downarrow$ Orb. $\downarrow$ Spec. $\downarrow$ Avg. $\downarrow$
GraphRNN (You et al., 2018) 0.0055 0.0584 0.0785 0.0065 0.0372
GRAN (Liao et al., 2019) 0.0113 0.0553 0.0540 0.0054 0.0315
EDP-GNN (Niu et al., 2020) 0.0011 0.0552 0.0520 0.0070 0.0288
SPECTRE (Martinkus et al., 2022) 0.0015 0.0521 0.0412 0.0056 0.0251
GDSS (Jo et al., 2022) 0.0212 0.0646 0.0894 0.0128 0.0470
DiGress (Vignac et al., 2023) 0.0013 0.0498 0.0434 0.0400 0.0336
GruM (Jo et al., 2024) 0.0015 0.0589 0.0450 0.0077 0.0283
DeFoG (Qin et al., 2025) 0.0006 0.0517 0.0556 0.0054 0.0283
HOG-Diff (Ours) 0.0028 0.0500 0.0428 0.0043 0.0249
F.3 Rationale for Selecting Topological Filters
@@
Overall, the proposed framework is most effective when the intermediate skeletons reflect structures naturally present in the data. This aligns with prior work in graph representation learning, which shows that the benefits of higher-order representations are most pronounced in datasets rich in such structures (Huang et al., 2024). Indeed, many real-world networks—including social, biological, and citation networks—exhibit abundant higher-order cells/simplices. Thus, many real-world tasks naturally fall into the “topologically rich” regime where our method is particularly advantageous.
Table 8: Average counts of higher-order structures across datasets. “–” indicates computation exceeded 10h.
Dataset 2-simplices 3-simplices 4-simplices 2-cells 3-cells
Community-small 27.45 8.24 0.76 40.92 0.00
Ego-small 4.61 1.64 0.35 5.66 0.00
Enzymes 25.93 3.05 0.01
SBM 441.25 90.88 4.42
QM9 0.47 0.00 0.00 1.84 0.10
ZINC250k 0.06 0.00 0.00 2.77 0.00
F.4 De Novo Molecule Generation
To visually assess the capability of the proposed method in molecular generation, Fig. 8 compares distributional results on the molecular datasets. Specifically, we first calculate the Morgan fingerprints (Rogers & Hahn, 2010) of all molecules, which have been widely utilized in drug discovery for capturing structural information. Subsequently, we apply Uniform Manifold Approximation and Projection (UMAP), a nonlinear dimensionality-reduction method that preserves local similarities, to embed the fingerprints into two dimensions and plot the resulting distributions.
We provide the standard deviation results and the additional validity (Val.) metric in Tabs. 9 and 10. Baseline results are sourced from Jo et al. (2022); Kong et al. (2023) or reproduced using the corresponding publicly available code.
Table 9: Comparison of different methods on QM9. We report the means and standard deviations of 3 runs. Asterisks (*) indicate that the source did not report standard deviations. The best results for the first three metrics are highlighted in bold.
[…]
Val. w/o
[…]
GraphArm* 0.002 1.220 90.25 100.00 95.62 70.39
MiCaM* 0.001 1.045 99.93 100.00 93.89 83.25
[…]
SPECTRE* 0.163 47.960 87.30 100.00 35.70 97.28
GSDM* 0.003 2.650 99.90 100.00 - -
[…]
DiGress* 0.0005 0.360 99.00 100.00 96.66 33.40
[…]
CatFlow* - 0.441 99.81 100.00 99.95 -
[…]
Table 10: Comparison of different methods on ZINC250k. We report the means and standard deviations of 3 runs. Asterisks (*) indicate that the source did not report standard deviations. The best results for the first three metrics are highlighted in bold.
[…]
Val. w/o
[…]
GraphArm* 0.055 16.260 88.23 100.00 99.46 100.00
MiCaM* 0.166 31.495 100.00 100.00 88.48 99.98
[…]
SPECTRE* 0.109 18.440 90.20 100.00 67.05 100.00
GSDM* 0.017 12.956 92.70 100.00 - -
[…]
DiGress* 0.082 23.060 91.02 100.00 81.23 100.00
[…]
CatFlow* - 13.211 99.95 99.99 100.00 -
[…]
F.5 Visualization Results
In this section, we additionally provide the visualizations of the generated graphs for both molecule generation tasks and generic graph generation tasks. Figs. 9-14 illustrate non-curated generated samples. HOG-Diff demonstrates the capability to generate high-quality samples that closely resemble the topological properties of empirical data while preserving essential structural details.
@@ F.5 Visualization Results
Appendix G Limitation
@@
As discussed in Sec. 4.2, the performance of the proposed framework depends on the presence of explicit higher-order structures. While previous studies have shown the prevalence of such structures in empirical systems, certain types of graphs, such as the ego-small dataset, lack this topological richness. In these cases, the benefits of higher-order diffusion guidance diminish, and the performance advantage becomes less pronounced.

### v3

@@ E.1 Computing Resources
E.2 Overview of Compared Methods
For benchmarking, we include various representative molecular generation models. Autoregressive models include GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021), and GraphArm (Kong et al., 2023); the fragment-based MiCaM (Geng et al., 2023) creatively leverages motif information. In contrast, the remaining methods adopt a one-shot generation paradigm: GraphEBM serves as an energy-based model, SPECTRE (Martinkus et al., 2022) and GSDM (Luo et al., 2023a) incorporate spectral conditioning within GAN and diffusion frameworks, respectively, while EDP-GNN (Niu et al., 2020), GDSS (Jo et al., 2022), DiGress (Vignac et al., 2023), and Cometh (Siraudin et al., 2025) represent diffusion-based generation models. We also compare against advanced flow-based models, including MoFlow (Zang and Wang, 2020), CatFlow (Eijkelboom et al., 2024), and DeFoG (Qin et al., 2025).
E.3 Molecule Generation
@@
We evaluate the quality of generated molecules on two well-known molecular datasets: QM9 (Ramakrishnan et al., 2014) and ZINC250k (Irwin et al., 2012), and obtain the intermediate higher-order skeletons using the 2-cell complex filtering. In experiments, each molecule is preprocessed into a graph comprising adjacency matrix $\bm{A}\in\{0,1,2,3\}^{n\times n}$ and node feature matrix $\bm{X}\in\{0,1\}^{n\times d}$, where $n$ denotes the maximum number of atoms in a molecule of the dataset, and $d$ is the number of possible atom types. The entries of $\bm{A}$ indicate the bond types: 0 for no bond, 1 for the single bond, 2 for the double bond, and 3 for the triple bond. Further, we scale $\bm{A}$ with a constant scale of 3 in order to bound the input of the model in the interval [0, 1], and rescale the final sample of the generation process to recover the bond types. Following the standard procedure (Shi et al., 2020; Luo et al., 2021), all molecules are kekulized by the RDKit library (Landrum and others, 2016) with hydrogen atoms removed. In addition, we make use of the valency correction proposed by Zang and Wang (2020). After generating samples by simulating the reverse diffusion process, the adjacency matrix entries are quantized to discrete values 0,1,2,3{0,1,2,3} by applying value clipping. Specifically, values in $(-\infty,0.5)$ are mapped to 0, $[0.5,1.5)$ to 1, $[1.5,2.5)$ to 2, and $2.5,+\infty)$ to 3, ensuring the bond types align with their respective categories.
We report the baseline results taken from Jo et al. (2022) and Kong et al. (2023) or as reproduced using publicly available code.
To comprehensively assess the quality of the generated molecules across datasets, we evaluate 10,000 generated samples using several key metrics: validity, validity with correction (Val. w/ corr.), uniqueness, novelty (Jo et al., 2022), Frechet ChemNet Distance (FCD) (Preuer et al., 2018), and Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa and Grave, 2010).
  • Validity computes the fraction of valid molecules before any corrections Jo et al. (2022), providing insight into the intrinsic quality of the generative process. While validity with correction (Val. w/ corr.) is measured as the fraction of valid molecules to all generated molecules after applying post-processing corrections such as valency adjustments or edge resampling. Whether molecules are valid is generally determined by compliance with the valence rules in RDKit (Landrum and others, 2016).
  • Novelty assesses the model’s ability to generalize by calculating the percentage of generated graphs that are not subgraphs of the training set, with two graphs considered identical if they are isomorphic.
  • Uniqueness quantifies the diversity of generated molecules as the ratio of unique samples to valid samples, removing duplicates that are subgraph-isomorphic, ensuring variety in the output.
  • FCD quantifies the similarity between generated and test molecules by leveraging the activations of ChemNet’s penultimate layer, assessing the generation quality within the chemical space.
  • NSPDK-MMD evaluates the generation quality from the graph topology perspective by computing the MMD between the generated and test sets while considering both node and edge features.
MOSES
We conducted experiments on the MOSES dataset (Polykovskiy et al., 2020), a refinement of the ZINC database for molecular generation containing approximately 1.9 million lead-like molecules. This dataset serves as a rigorous testbed for assessing whether the proposed higher-order guidance remains computationally feasible and effective at scale. Following Vignac et al. (2023), the reported scores for FCD, SNN, and Scaffold similarity are computed on the dataset made of separate scaffolds, which measures the ability of the networks to predict new ring structures.
GuacaMol
GuacaMol (Brown et al., 2019) is derived from the ChEMBL database and contains 1.4M molecules, from which 1.1M are used for training. We apply a preprocessing step similar to Vignac et al. (2023), filtering out molecules that cannot be mapped from SMILES to a graph and back to SMILES. In Fig. 3(left), we report the scores of the FCD and KL metrics following Brown et al. (2019). Note that, unlike other settings, the FCD score is normalized using $FCDscore=\exp(-0.2\cdot FCD)$ to obtain a final value between 0 and 1, meaning that higher values indicate better performance.
For the KL score, we compute the following descriptors for both the generated and reference molecules: BertzCT, MolLogP, MolWt, TPSA, NumHAcceptors, NumHDonors, NumRotatableBonds, NumAliphaticRings, NumAromaticRings, and the ECFP4 fingerprint-based similarity to the nearest neighbor. The KL divergence $D_{KL,i}$ is computed for each descriptor to measure the difference between the two molecular sets. These divergences are then combined to produce a final normalized score $S=\frac{1}{k}\sum_{i=1}^{k}\exp(-D_{KL,i})$.
E.4 Generic Graph Generation
@@
To display the topology distribution learning ability, we assess HOG-Diff over four common generic graph datasets: (1) Community-small, containing 100 randomly generated community graphs; (2) Ego-small, comprising 200 small ego graphs derived from the Citeseer network dataset; (3) Enzymes, featuring 587 protein graphs representing tertiary structures of enzymes from the BRENDA database; and (4) Stochastic Block Model (SBM), a larger-scale dataset.
Community-small, Ego-small, and Enzymes
@@
SBMs
To further evaluate the scalability and robustness of HOG-Diff, we conduct experiments on the Stochastic Block Model (SBM) dataset, which comprises graphs of larger scale and is evaluated under a distinct experimental protocol from that used in Sec. 4.2. The dataset consists of 200 synthetic graphs generated using the stochastic block model. The number of communities is uniformly sampled between 2 and 5, and the number of nodes within each community is uniformly sampled between 20 and 40. Edges are created with probabilities of 0.3 for intra-community connections and 0.05 for inter-community connections.
We adopt the evaluation setting introduced by Martinkus et al. (2022), including the same data splits, feature initialization, and the use of Total Variation (TV) distance to compute the MMDs. The TV kernel is adopted since it offers higher computational efficiency compared to the Earth Mover’s Distance (EMD) kernel, especially for large graphs. We also report the percentage of valid, unique, and novel (V.U.N.) samples among the generated graphs to further assess the ability of our model to capture the properties of the targeted distributions correctly. Baseline results are sourced from Vignac et al. (2023); Jo et al. (2024) or reproduced using the corresponding publicly available code.
E.5 Data Statistics
Tab. 7 summarizes the key characteristics of the datasets utilized in this study. The table outlines the type of dataset, the total number of graphs, and the range of graph sizes ($|V|$). Additionally, it also provides the number of distinct node types and edge types for each dataset. Notably, the Community-small and Ego-small datasets contain relatively small graphs, whereas Enzymes, SBM and the molecular datasets exhibit greater diversity in terms of graph size and complexity.
@@ E.5 Data Statistics
Dataset Graph type #Graphs #Nodes Node types Edge types
@@ Appendix F Additional Experimental Results
F.1 Extended Results on Standard Molecular Benchmarks: QM9 and ZINC250k
To visually assess the capability of the proposed method in molecular generation, Fig. 9 compares distributional results on the molecular datasets. Specifically, we first calculate the Morgan fingerprints (Rogers and Hahn, 2010) of all molecules, which have been widely utilized in drug discovery for capturing structural information. Subsequently, we apply Uniform Manifold Approximation and Projection (UMAP), a nonlinear dimensionality-reduction method that preserves local similarities, to embed the fingerprints into two dimensions and plot the resulting distributions.
We provide the standard deviation results and the additional metric validity with correction (Val. w/ corr.) in Tabs. 8 and 9. Baseline results are sourced from Jo et al. (2022); Kong et al. (2023) or reproduced using the corresponding publicly available code.
Table 8: Comparison of different methods on QM9. We report the means and standard deviations of 3 runs. The best results are highlighted in bold. Asterisks (*) indicate that the source did not report standard deviations.
[…]
GraphArm* (Kong et al., 2023) 90.25 100.00 95.62 70.39 1.220 0.002
MiCaM* (Geng et al., 2023) 99.93 100.00 93.89 83.25 1.045 0.001
[…]
SPECTRE* (Martinkus et al., 2022) 87.30 100.00 35.70 97.28 47.960 0.163
GSDM* (Luo et al., 2023a) 99.90 100.00 - - 2.650 0.003
[…]
DiGress* (Vignac et al., 2023) 99.00 100.00 96.66 33.40 0.360 0.0005
[…]
CatFlow* (Eijkelboom et al., 2024) 99.81 100.00 99.95 - 0.441 -
[…]
Table 9: Comparison of different methods on ZINC250k. We report the means and standard deviations of 3 runs. The best results are highlighted in bold. Asterisks (*) indicate that the source did not report standard deviations.
[…]
GraphArm* (Kong et al., 2023) 88.23 100.00 99.46 100.00 16.260 0.055
MiCaM* (Geng et al., 2023) 100.00 100.00 88.48 99.98 31.495 0.166
[…]
SPECTRE* (Martinkus et al., 2022) 90.20 100.00 67.05 100.00 18.440 0.109
GSDM* (Luo et al., 2023a) 92.70 100.00 - - 12.956 0.017
[…]
DiGress* (Vignac et al., 2023) 91.02 100.00 81.23 100.00 23.060 0.082
[…]
CatFlow* (Eijkelboom et al., 2024) 99.95 99.99 100.00 - 13.211 -
[…]
F.2 Scalability Evaluation on Large Benchmarks: MOSES and SBM
To address concerns regarding scalability and performance on larger benchmarks, we extended our evaluation to encompass both large-scale molecular datasets and generic graph benchmarks.
MOSES
We conducted experiments on the MOSES dataset (Polykovskiy et al., 2020). The results are presented in Tab. 10. HOG-Diff demonstrates superior performance in terms of distribution learning, achieving the lowest FCD score. This indicates that our method more accurately captures the underlying chemical and topological distribution of the large-scale training set.
Table 10: Generation performance on the large-scale MOSES dataset. Top results are highlighted in bold.
Method Val. $\uparrow$ Uni. $\uparrow$ Nov. $\uparrow$ Filters $\uparrow$ FCD $\downarrow$ SNN Scaf
GraphINVENT (Mercado et al., 2021) 96.4 99.8 - 95.0 1.22 0.54 12.7
DiGress (Vignac et al., 2023) 85.7 100.0 95.0 97.1 1.19 0.52 14.8
DisCo (Xu et al., 2024) 88.3 100.0 97.7 95.6 1.44 0.50 15.1
Cometh (Siraudin et al., 2025) 90.5 99.9 92.6 99.1 1.27 0.54 16.0
DeFoG (Qin et al., 2025) 92.8 99.9 92.1 98.9 1.95 0.55 14.4
[…]
SBMs
To further evaluate the scalability and robustness of HOG-Diff, we report results on the Stochastic Block Model (SBM) dataset in Tab. 11, which comprises graphs of larger scale and is evaluated under a distinct experimental protocol from that used in Sec. 4.2. As shown in Tab. 11, HOG-Diff achieves competitive performance compared to the state-of-the-art.
[…]
GraphRNN (You et al., 2018) 0.0055 0.0584 0.0785 0.0065 0.0372 0.05
GRAN (Liao et al., 2019) 0.0113 0.0553 0.0540 0.0054 0.0315 0.25
EDP-GNN (Niu et al., 2020) 0.0011 0.0552 0.0520 0.0070 0.0288 -
SPECTRE (Martinkus et al., 2022) 0.0015 0.0521 0.0412 0.0056 0.0251 0.53
GDSS (Jo et al., 2022) 0.0212 0.0646 0.0894 0.0128 0.0470 0.05
RefineGen (Bergmeister et al., 2024) 0.0141 0.0528 0.0809 0.0071 0.0387 0.75
GruM (Jo et al., 2024) 0.0015 0.0589 0.0450 0.0077 0.0283 0.85
DeFoG (Qin et al., 2025) 0.0006 0.0517 0.0556 0.0054 0.0283 0.80
HOG-Diff (Ours) 0.0028 0.0500 0.0428 0.0043 0.0249 0.83
F.3 Analysis of Diffusion Domain Choice
@@
F.4 Rationale for Selecting Topological Filters
@@
Overall, the proposed framework is most effective when the intermediate skeletons reflect structures naturally present in the data. This aligns with prior work in graph representation learning, which shows that the benefits of higher-order representations are most pronounced in datasets that are rich in such structures (Huang et al., 2024). Indeed, many real-world networks, including social, biological, and citation networks, exhibit abundant higher-order cells and simplices. Thus, many real-world tasks naturally fall into the “topologically rich” regime in which our method is particularly advantageous.
Table 13: Average counts of simplices and cells across datasets. “-” indicates computation exceeded 10h.
Dataset 0-simplices 1-simplices 2-simplices 3-simplices 4-simplices 2-cells 3-cells
Community-small 15.28 35.15 27.45 8.24 0.76 40.92 0.00
Ego-small 6.41 8.70 4.61 1.64 0.35 5.66 0.00
Enzymes 33.03 63.27 25.93 3.05 0.01
SBM 104.01 500.41 441.25 90.88 4.42
QM9 8.80 9.40 0.47 0.00 0.00 1.84 0.10
ZINC250k 23.15 24.90 0.06 0.00 0.00 2.77 0.00
F.5 Additional Visualizations
Fig. 10 provides a qualitative view of the generative trajectory of HOG-Diff for the model trained on ZINC250k. It shows several intermediate graph states from the initial noisy prior to the final molecule, illustrating the coarse-to-fine refinement process.
In this section, we additionally provide the visualizations of the generated graphs for both molecule generation tasks and generic graph generation tasks. Figs. 11-16 illustrate non-curated generated samples. HOG-Diff demonstrates the capability to generate high-quality samples that closely resemble the topological properties of empirical data while preserving essential structural details.
@@ F.5 Additional Visualizations
Appendix G Limitations and Further Work
@@
As discussed in Sec. 4.2, the performance of the proposed framework depends on the presence of explicit higher-order structures. Although previous studies have shown that such structures are prevalent in many empirical systems, certain types of graphs, such as the ego-small dataset, lack this topological richness. In these cases, the benefits of higher-order diffusion guidance diminish, and the performance advantage becomes less pronounced.